A hemispherical wash basin has an internal diameter of 22 cm. The basin has a uniform thickness of 10 cm. What is the area of the inner and outer surfaces of the wash basin? Answer in square centimetres correct to four significant figures.
Please show working out

Answer by rothauserc(4718)

The curved surface area of a hemisphere is given by the formula
SA = 2*pi*r^2
The radius of the inner surface is 22/2 = 11, so
SA of inner surface = 2*pi*11^2 = 760.26542216872996370796 = 760.2654 square cm
The radius of the outer surface is 11 + 10 = 21
SA of the outer surface = 2*pi*21^2 = 2770.88472046619763632405 = 2770.8847 square cm